Juan's band produced a number of CDs to sell at a gig. The batch of CDs cost $170 to make and each CD was sold for $2.50 more than it cost to produce. Juan gave 2 CDs to his friends, but sold the rest for a total of $198. How many CDs did Juan produce?

So i said that $5 in profit= 2 CDs. Then i did 198-170= $28 profit. Then I did 28/5 to get the amount of CDs he created. I think that was wrong because it came out as a decimal answer and you can't really have part of a CD. I don't know who to solve this.

Answer by mananth(16946)

Juan's band produced a number of CDs to sell at a gig. The batch of CDs cost $170 to make and each CD was sold for $2.50 more than it cost to produce. Juan gave 2 CDs to his friends, but sold the rest for a total of $198. How many CDs did Juan produce?
let him produce x CDs
cost of each CD =170/x
Selling price of each CD = 170/x +5/2
he sold 2 less ie:(x-2)
Total sales = 198 = (170/x +5/2)(x-2)
198= ((340+5x)/2x)(x-2)
396x=(340+5x)(x-2)
396x=340x-680+5x^2-10x
5x^2-66x-680=0
solve for x
5x^2-100x+34x-680=0
5x(x-20)+34(x-20)=0
(x-20)(5x+34)=0
x=20 taking the positive root
he produced 20 CDs